OpenStudy (anonymous):

suppose that f(x) = 4x+7 and g(x) =2x. what is a rule for g(f(x)? what is a rule for f(g(x))? suppose that f(x) = 4x+7 and g(x) =2x. what is a rule for g(f(x)? what is a rule for f(g(x))? @Mathematics

OpenStudy (amistre64):

g(f(x)) just means you plug in f(x) into the function of g everywhere you see an x in g same concept for f(g(x)) but in the opposite order

OpenStudy (amistre64):

g(x) =2x g(f(x)) = 2(f(x))

OpenStudy (amistre64):

f(x) = 4x+7 f(g(x)) = 4(g(x))+7

OpenStudy (anonymous):

oh ok thanks so much

OpenStudy (amistre64):

yep